OpenStudy (anonymous):

Refer to the graph of y = sin x or y = cos x to find the exact values of x in the interval [0, 4π] that satisfy the equation. (Enter your answers as a comma-separated list.) 5 sin x = 5/2

OpenStudy (anonymous):

With a little algebra \[\sin x=\frac{1}{2}\] so... where on the interval from 0 to 4pi does the sin x = (1/2)?

OpenStudy (anonymous):

umm pi/3?

OpenStudy (anonymous):

Hint... the arcsin(1/2)= (pi/6) there are 4 spots